the air in a 5.000-L tank has a pressure of 1.20 atm. What is the final pressure, in atmospheres, when the air is placed in tanks that have the following volumes, if there is not change in temperature and amount of gas? a) 1.00L b) 2500.mL c)750.mL d) 8.00L
Answer #1

a.

   intial                                    final

P1 = 1.2atm                          P2 =

V1   = 5L                                  V2 = 1L

           P1V1   =   P2V2

               P2   = P1V1/V2

                        = 1.2*5/1   = 6atm>>>>>>>>>answer

b.

intial                                    final

P1 = 1.2atm                          P2 =

V1   = 5L                                  V2 = 2500ml   = 2.5L

           P1V1   =   P2V2

               P2   = P1V1/V2

                       = 1.2*5/2.5   = 2.4atm

c.

intial                                    final

P1 = 1.2atm                          P2 =

V1   = 5L                                  V2 = 750ml   = 0.75L

           P1V1   =   P2V2

               P2   = P1V1/V2

                       = 1.2*5/0.75 = 8atm

d.

intial                                    final

P1 = 1.2atm                          P2 =

V1   = 5L                                  V2 = 8L

           P1V1   =   P2V2

               P2   = P1V1/V2

                       = 1.2*5/8 = 0.75atm